# Building-block flow model for computational fluids

### v1

@@ Abstract
We introduce a closure model for computational fluid dynamics, referred to as the Building-block Flow Model (BFM). The foundation of the model rests on the premise that a finite collection of simple flows encapsulates the essential physics necessary to predict more complex scenarios. The BFM is implemented using artificial neural networks and introduces five unique advancements within the framework of large-eddy simulation: (1) It is designed to predict multiple flow regimes (wall turbulence under zero, favorable, adverse mean-pressure-gradient, and separation); (2) It unifies the closure model at solid boundaries (i.e., the wall model) and the rest of the flow (i.e., the subgrid-scale model) into a single entity; (3) It ensures consistency with numerical schemes and gridding strategy by accounting for numerical errors; (4) It is directly applicable to arbitrary complex geometries; (5) It can be scaled up to model additional flow physics in the future if needed (e.g., shockwaves and laminar-to-turbulent transition). The BFM is utilized to predict key quantities of interest in turbulent channel flows, a Gaussian bump, and an aircraft in a landing configuration. In all cases, the BFM demonstrates similar or superior capabilities in terms of accuracy and computational efficiency compared to previous state-of-the-art closure models.
@@ 1 Introduction
In terms of prediction and control, the field of CFD is in the privileged position of having a highly accurate framework to model flows: the Navier–Stokes equations. However, the number of degrees of freedom involved in real-world applications ranges between $10^{13}$ to $10^{20}$, making the direct numerical simulation of all flow scales infeasible. Consequently, the treatment of turbulence in industrial CFD has primarily relied on closure models for the Reynolds-averaged Navier-Stokes (RANS) equations [7]. The approach has materialized in various forms, from pure RANS solutions to hybrid methods such as Detached Eddy Simulation and its variants [8]. Many RANS models have been developed to overcome the limitations of their predecessors, often by expanding and calibrating their coefficients to account for missing physics. However, no practical model is universally applicable across the broad range of flow regimes of interest to the industry. Examples of challenging flow scenarios include separated flows, favorable/adverse pressure gradient effects, shock waves, mean-flow three-dimensionality, and transition to turbulence, among others, as illustrated in Fig. 1(left). In these situations, RANS predictions tend to be inconsistent and unreliable, especially for geometries and conditions representative of the flight envelope of commercial airplanes [9]. Further CFD experience with aircraft at high angles of attack has revealed that RANS-based solvers have difficulties predicting maximum lift and the physical mechanisms for stall [10]. As a result, the number of wind tunnel experiments required during the final stage of the aircraft design cycle has remained at around ten for the past two decades (Fig. 1 right).
@@
Figure 1: (Left) Schematic of the different flow phenomena encountered over an aircraft challenging current CFD methodologies. For compactness, all the phenomena are overlaid in the schematic; however, not all of them occur concurrently or even take place at all. (Right) Number of wind tunnel tests required during the design cycle as a function of time. The horizontal axis also marks the introduction of widely used RANS models. The graph is adapted from ref.[11].
@@
Recently, large-eddy simulation (LES) has gained momentum as a tool for both scientific investigations and industrial applications. In LES, the large flow motions containing most of the energy are directly resolved by the grid, while the dissipative effect of the small scales is modeled through a subgrid-scale (SGS) model. By additionally modeling the near-wall flow using a so-called wall model, the grid-point requirements for wall-modeled LES (WMLES) scale at most linearly with increasing Reynolds number [12, 13, 14], defined as the ratio between the largest and smallest flow scales in the problem. Most widely used SGS models were developed from the 1980s to the early 2000s and are based on the eddy-viscosity assumption [15]. Among the most prominent SGS models, we can cite the similarity model [16], Vreman model [17], dynamic Smagorinsky model [18], deconvolution model [19], and sigma model [20], to name a few. A detailed account of SGS models can be found in ref.[21]. Regarding wall modeling, several strategies have been explored in the literature, and comprehensive reviews can be found in refs. [22, 23, 8, 24, 25]. A widely adopted method for wall modeling is the wall-flux approach. This approach substitutes the no-slip and thermal boundary conditions at the wall with the shear stress and heat flux values provided by the wall model. Popular examples of these approaches include those based on the law of the wall [26, 27, 28], the full/simplified RANS equations [29, 30, 31, 32, 33, 34, 35], vortex-based models [36], or dynamic wall models [37, 38].
@@
In this context, machine learning (ML) has emerged as a powerful tool to enhance existing turbulence modeling approaches in the fluid community[39, 40, 41, 42, 43, 44]. Over the last two decades, there have been multiple efforts devoted to the development of SGS and wall models using ML tools. Most ML-based SGS models rely on supervised learning, which involves the discovery of a function that maps an input to an output based on provided training input-output pairs. Early approaches used artificial neural networks (ANNs) to emulate and speed up conventional, yet computationally expensive, SGS models[45]. More recently, SGS models have been trained to predict the so-called perfect SGS terms using data from filtered direct numerical simulation (DNS)[46, 47, 48]. Other approaches involve deriving SGS terms from optimal estimator theory[49] and deconvolution operators[50, 51, 52], as well as from reinforcement learning[53]. Similar ML methods have been employed to devise wall models for LES via supervised learning. One of the initial attempts can be found in ref. [54]. Other examples include spanwise rotating channel flows[55], flows over periodic hills[56], turbulent flows with separation[57], and boundary layer flows in the presence of shock-boundary layer interaction[58], with mixed results in a-posteriori testing. Recent works have also leveraged semi-supervised learning to devise wall models[59], although these approaches are still limited to simple flow configurations.
@@
2 Results
2.1 The building-block flow model
@@
2.2 Building block flows
The BBFs offer simple representations of different flow regimes for which scale-resolving, high-fidelity DNS data is available. The configuration of the BBFs entails an incompressible turbulent flow confined between two parallel walls. Four types of BBFs are considered: wall-bounded turbulence under separation, adverse, zero, and favorable mean pressure gradient. Examples of these BBFs are illustrated in Fig. 3. The intentional exclusion of cases with additional complexities, such as airfoils, wings, bumps, etc., is grounded in the idea that the BBFs should encapsulate fundamental flow physics to predict more intricate scenarios. Hence, the aim is to avoid case overfitting, e.g., correctly predicting the flow over a wing merely because the model was trained on similar wings. Note that the largest scales of the flow are intended to be resolved by the computational grid, whereas the BBFs represent the smaller flow features. This justifies the geometric simplicity of the BBFs over more complex geometries. Additionally, the computational affordability of the BBFs allows for the collection of a large number of high-fidelity simulations for training, which facilitates efficient exploration of multiple flow regimes. This aspect is important as the BFM must learn the physical scaling of the non-dimensional inputs and outputs that control the problem at hand, rather than merely fitting to a set of cases.
@@ 2.3 Model architecture
2.4 Training data
@@
2.5 Validation cases
We validate the performance of the BFM in three cases: a turbulent channel flow, the separated flow over a bump, and an aircraft in landing configuration. The results are compared with simulations using established SGS and wall models conducted using identical meshes as those for the BFM. The models for comparison are the dynamic Smagorisky SGS model (DSM) [18, 62] or Vreman SGS model (VRE) [17] combined with the equilibrium wall model (EQ) [33]. The models are labeled as DSM-EQ and VRE-EQ, respectively. In the validation cases below, the computational cost of conducting WMLES with BFM is approximately 0.9 times that of DSM-EQ and 1.1 times that of VRE-EQ.
2.6 Turbulent channel flow
@@
2.7 Separated flow over a bump
@@
Figure 5: (A) Schematic of the Gaussian bump; (B) grid structure and (C) instantaneous streamwise velocity in the $y/L=0$ plane. The velocity of the freestream is $U_{\infty}$. Note that the figure does not show the whole domain in the streamwise direction. (D) Wall pressure coefficient $C_{p}=\frac{p-p_{\infty}}{\frac{1}{2}\rho_{\infty}U^{2}_{\infty}}$ and (E) wall friction coefficient $C_{f}=\frac{\tau_{w}}{\frac{1}{2}\rho_{\infty}U^{2}_{\infty}}$, along the $y/L=0$ plane, where $U_{\infty}$, $p_{\infty}$ and $\rho_{\infty}$ are the freestream velocity, pressure and density, respectively, and $p$ and $\tau_{w}$ are averaged over time. (F) Mean velocity profile at $y/L=0$ at different streamwise locations. In (D,E,F) The solid lines denote BFM (blue), DSM-EQ (green) and VRE-EQ (orange). White circles are experimental values[74].
@@
The second validation case is a Gaussian bump designed by The Boeing Company and the University of Washington [75], as shown in Fig. 5. The test consists of a three-dimensional tapered hump at a Reynolds number of $3.6\times 10^{6}$ and a Mach number of $0.17$, based on the freestream quantities and the length of the bump $L$. This test case has proven to be highly challenging for the CFD community, with inaccurate prediction of the separation bubble and non-monotonic convergence of the solution with grid refinements[76, 77].
@@
We evaluate the performance of the BFM in predicting the mean velocity profiles, and the mean pressure and friction at the wall (characterized by the non-dimensional coefficients $C_{p}$ and $C_{f}$, respectively). The simulation consists of roughly 9 million control volumes. Similar WMLESes are conducted using DSM-EQ and VRE-EQ. The numerical findings are compared with experimental data[74]. Figure 5(D,E) illustrates the evolution of $C_{p}$ and $C_{f}$ along the centerline of the bump ($x/L$). Notably, the BFM yields more accurate predictions of the $C_{p}$ and $C_{f}$ evolution in the separation region ($x/L>0.1$) compared to DSM-EQ and VRE-EQ, which even fail to predict flow separation. The improved predictions by the BFM also extend to the mean velocity profiles, as demonstrated in Fig. 5F. The BFM predicts mean velocity with up to 3% accuracy, whereas both DSM-EQ and VRE-EQ yield a similar result that overestimate the mean velocities by more than 30%. Nevertheless, there is still room for improvement upstream of the bump ($x/L<0$), where all models exhibit low accuracy in the prediction of $C_{f}$.
2.8 Aircraft in landing configuration
@@
The third validation case is an aircraft in landing configuration, specifically the NASA Common Research Model High-Lift (CRM-HL). This case serves as the gold standard used by the aerospace community to assess the capabilities of different CFD methodologies [78]. The CRM-HL is a geometrically realistic aircraft that includes the bracketry associated with deployed flaps and slats, as well as a flow-through nacelle mounted on the underside of the wing (Fig. 2). The Reynolds number of the aircraft is 5.49 million based on the mean aerodynamic chord and freestream velocity, and the freestream Mach number is 0.20.
@@
The simulations are conducted using approximately 30 million control volumes. Two angles of attack are considered: $\alpha=7.05^{\circ}$ and $19.57^{\circ}$. Figure 2 contains the model geometry, an inset of the mesh, and a visualization of the wall-shear stress predicted by BFM.
@@
Figure 6 displays the forces acting on the aircraft, characterized by the lift ($C_{L}$), drag ($C_{D}$), and pitching moment ($C_{M}$) coefficients for BFM, DSM-EQ, and VRE-EQ, compared with experimental data [79]. For $\alpha=7.05^{\circ}$, DSM-EQ and VRE-EQ predict values of $C_{L}$ and $C_{D}$ close to experimental results. However, further analysis has shown that this accurate prediction is coincidental and due to error cancellation during the integration of total forces [80]. This is corroborated by the poorer predictions of the pitching moment by DSM-EQ and VRE-EQ. The case of $\alpha=19.57^{\circ}$ corresponds to the angle of attack of the maximum lift coefficient. DSM-EQ and VRE-EQ underpredict $C_{L}$, whereas the BFM offers improved predictions. Inspection of the sectional pressure coefficient, shown in Fig. 7, indicates that the BFM maintains the flow attached over a longer section of the wing compared to DSM-EQ and VRE-EQ, explaining the enhanced performance of the BFM. Overall, BFM yields improved predictions for the three coefficients compared to traditional models. The results are still far from satisfactory, especially for the pitching moment, and there is clear potential for further improvements. Nonetheless, it is worth noting that BFM has never ‘seen’ an aircraft-like flow or been trained in a case resembling an airfoil or a wing. This demonstrates the ability of the BFM methodology to offer predictions that go beyond simple regression.
Figure 6: (A) Lift coefficient, (B) drag polar and (C) pitching moment coefficient for the CRM-HL. The black lines denote experimental results, blue circles are for the BFM, squares are for DSM-EQ and orange squared for VRE-EQ.
@@
Figure 7: Pressure coefficient along the chord of the wing at the 82% spanwise section of the wing for $\alpha=19.57^{\circ}$. The chord location is normalized by the sectional chord-length $c_{\text{ref}}$. White circles represents experimental results, blue circles are for the BFM, orange and green squares are for DSM-EQ and VRE-EQ, respectively.
3 Discussion
3.1 Opportunities of BFM for CFD

### v2

# Building-block flow model for large-eddy simulation

@@ Abstract
We introduce a closure model for wall-modeled large-eddy simulation (WMLES), referred to as the Building-block Flow Model (BFM). The foundation of the model rests on the premise that a finite collection of simple flows encapsulates the essential physics necessary to predict more complex scenarios. The BFM is implemented using artificial neural networks and introduces five advancements within the framework of WMLES: (1) It is designed to predict multiple flow regimes (wall turbulence under zero, favorable, adverse mean-pressure-gradient, and separation); (2) It unifies the closure model at solid boundaries (i.e., the wall model) and the rest of the flow (i.e., the subgrid-scale model) into a single entity; (3) It ensures consistency with numerical schemes and gridding strategy by accounting for numerical errors; (4) It is directly applicable to arbitrary complex geometries; (5) It can be scaled up to model additional flow physics in the future if needed (e.g., shockwaves and laminar-to-turbulent transition). The BFM is utilized to predict key quantities of interest in turbulent channel and pipe flows, a Gaussian bump, a simplified aircraft, and a realistic aircraft in landing configuration. In all cases, the BFM demonstrates similar or superior capabilities in terms of accuracy and computational efficiency compared to previous state-of-the-art closure models.
@@ 1 Introduction
In terms of prediction and control, the field of CFD is in the privileged position of having a highly accurate framework to model flows: the Navier–Stokes equations. However, the number of degrees of freedom involved in real-world applications ranges between $10^{13}$ to $10^{20}$, making the direct numerical simulation of all flow scales infeasible. Consequently, the treatment of turbulence in industrial CFD has primarily relied on closure models for the Reynolds-averaged Navier-Stokes (RANS) equations [7]. The approach has materialized in various forms, from pure RANS solutions to hybrid methods such as Detached Eddy Simulation and its variants [8]. Many RANS models have been developed to overcome the limitations of their predecessors, often by expanding and calibrating their coefficients to account for missing physics. However, no practical model is universally applicable across the broad range of flow regimes of interest to the industry. Examples of challenging flow scenarios include separated flows, favorable/adverse pressure gradient effects, shock waves, mean-flow three-dimensionality, and transition to turbulence, among others, as illustrated in Fig. 1(left). In these situations, RANS predictions tend to be inconsistent and unreliable, especially for geometries and conditions representative of the flight envelope of commercial airplanes [9]. Further CFD experience with aircraft at high angles of attack has revealed that RANS-based solvers have difficulties predicting maximum lift and the physical mechanisms for stall [10]. As a result, the number of wind tunnel experiments required during the final stage of the aircraft design cycle has remained at around ten for the past two decades (Fig. 1 top right).
Refer to captionRefer to captionSelected milestones in WMLES196519701975198019851990199520002005201020152020Smag.[11]LoW-WM[12]SSM[13]DSM[14]WALE[15], Deconv.-SGS[16]TBL-WM[17]ANN SGS[18]Vreman[19]EQ-WM[20]ANN WM[21] WMLES of an aircraft[22]
@@
Figure 1: (Left) Schematic of the different flow phenomena encountered over an aircraft challenging current CFD methodologies. For compactness, all the phenomena are overlaid in the schematic; however, not all of them occur concurrently or even take place at all. (Top right) Number of wind tunnel tests required during the design cycle as a function of time. The horizontal axis also marks the introduction of widely used RANS models. The graph is adapted from [23]. (Bottom right) Selected milestones in wall and SGS modeling for (WM)LES of external aerodynamics. The text in blue highlights milestones in traditional (WM)LES; the text in red highlights machine-learning based milestones in (WM)LES. In the figure, LoW stands for ‘law of the wall’ and TBL for ‘thin boundary layer’.
@@
Recently, large-eddy simulation (LES) has gained momentum as a tool for both scientific investigations and industrial applications. In LES, the large flow motions containing most of the energy are directly resolved by the grid, while the dissipative effect of the small scales is modeled through a subgrid-scale (SGS) model. By additionally modeling the near-wall flow using a so-called wall model, the grid-point requirements for wall-modeled LES (WMLES) scale at most linearly with increasing Reynolds number [24, 25, 26], defined as the ratio between the largest and smallest flow scales in the problem. Most widely used SGS models were developed from the 1980s to the early 2000s and are based on the eddy-viscosity assumption [11] (see Fig. 1 bottom right). Among the most prominent SGS models, we can cite the similarity model [13], Vreman model [19], dynamic Smagorinsky model [14], deconvolution model [16], and sigma model [27], to name a few. A detailed account of SGS models can be found in [28]. Regarding wall modeling, several strategies have been explored in the literature, and comprehensive reviews can be found in refs. [17, 29, 8, 30, 31]. A widely adopted method for wall modeling is the wall-flux approach. This approach substitutes the no-slip and thermal boundary conditions at the wall with the shear stress and heat flux values provided by the wall model. Popular examples of these approaches include those based on the law of the wall [12, 32, 33], the full/simplified RANS equations [34, 35, 36, 37, 20, 38, 39], vortex-based models [40], or dynamic wall models [41, 42].
@@
In this context, machine learning (ML) has emerged as a powerful tool to enhance existing turbulence modeling approaches in the fluid community[43, 44, 45, 46, 47, 48]. Over the last two decades, there have been multiple efforts devoted to the development of SGS and wall models using ML tools. Most ML-based SGS models rely on supervised learning, which involves the discovery of a function that maps an input to an output based on provided training input-output pairs. Early approaches used artificial neural networks (ANNs) to emulate and speed up conventional, yet computationally expensive, SGS models[18]. More recently, SGS models have been trained to predict the so-called perfect SGS terms using data from filtered direct numerical simulation (DNS)[49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Other approaches involve deriving SGS terms from optimal estimator theory[59] and deconvolution operators[60, 61, 62, 63, 64], as well as from reinforcement learning[65, 66, 67]. Similar ML methods have been employed to devise wall models for LES via supervised learning. One of the initial attempts can be found in ref. [21]. Other examples include spanwise rotating channel flows[68], flows over periodic hills[69], turbulent flows with separation[70], and boundary layer flows in the presence of shock-boundary layer interaction[71], with mixed results in a posteriori testing. Recent works have also leveraged semi-supervised learning to devise wall models[72], although these approaches are still limited to simple flow configurations.
@@
Results
The building-block flow model
@@ Results
Building block flows
The BBFs offer simple representations of different flow regimes for which scale-resolving, high-fidelity DNS data or experimental data are available. The configuration of the BBFs entails an incompressible turbulent flow confined between two parallel walls. Four types of BBFs are considered: wall-bounded turbulence under separation, adverse, zero, and favorable mean pressure gradient. Examples of these BBFs are illustrated in Fig. 3. The intentional exclusion of cases with additional complexities, such as airfoils, wings, bumps, etc., is grounded in the idea that the BBFs should encapsulate fundamental flow physics to predict more intricate scenarios. Hence, the aim is to avoid case overfitting, e.g., correctly predicting the flow over a wing merely because the model was trained on similar wings. Note that the largest scales of the flow are intended to be resolved by the computational grid, whereas the BBFs represent the smaller flow features. This justifies the geometric simplicity of the BBFs over more complex cases. Additionally, the computational affordability of the BBFs allows for the collection of a large number of high-fidelity data for training, which facilitates efficient exploration of multiple flow regimes. This aspect is important as the BFM must learn the physical scaling of the non-dimensional inputs and outputs that control the problem at hand, rather than merely fitting to a set of cases.
@@
Numerically-consistent training data
@@
Validation cases
We evaluate the performance of the BFM in five cases: turbulent channel flows, a turbulent pipe flow, the separated flow over a bump, a simplified aircraft model and a realistic aircraft in landing configuration. The results are compared with simulations using established SGS and wall models conducted using identical meshes as those for the BFM. The models for comparison are the dynamic Smagorisky SGS model (DSM) [14, 80] or Vreman SGS model (VRE) [19] combined with the equilibrium wall model (EQ) [20]. The models are labeled as DSM-EQ and VRE-EQ, respectively. In the validation cases below, the computational cost of conducting WMLES with BFM is approximately 0.9 times that of DSM-EQ and 1.1 times that of VRE-EQ.
Turbulent channel flow
@@
Turbulent pipe flow
The second validation case is the turbulent flow in a pipe (see Fig. 5A). We use this case to evaluate the performance of BFM at Reynolds numbers much higher than those it was trained for. The friction Reynolds number is $\textit{Re}_{\tau}=u_{\tau}R/\nu=39,500$, where $R$ is the pipe radius, and the flow is incompressible. The pipe is assumed periodic in the streamwise direction with a length of $7.5R$. As in the turbulent channel flow, the pipe is driven by a streamwise mean pressure gradient such that $\textit{Re}_{\textit{c}}=U_{\textit{c}}R/\nu\approx 1,272,800$, with $U_{\textit{c}}$ representing the mean streamwise velocity at the pipe centerline. The mesh is spatially isotropic (see Fig. 5A) with a grid size of $\Delta=0.1R$, which yields roughly 30,000 grid points.
We evaluate the performance of BFM, DSM-EQ, and VRE-EQ using the same flow conditions and computational grid. The results are compared with experimental data[85]. Figure 5B displays the mean velocity profile, showing the improved agreement of BFM with the experimental results with respect to DSM-EQ and VRE-EQ. Similarly, BFM predicts the average wall shear stress within a 3% error compared to the experiment, whereas DSM-EQ and VRE-EQ overpredict the average wall shear stress by 22% and 13%, respectively.
Figure 5: (A) Schematic and flow visualization of the pipe flow at $\textit{Re}_{\tau}=39500$. The radial and axial slices depict the grid structure and the instantaneous streamwise velocity. (B) Mean velocity profile along the radial direction $r$ from the wall ($r/R=0$) to the pipe centerline ($r/R=1$). Symbols correspond to BFM (blue circles), DSM-EQ (red squares), and VRE-EQ (yellow triangles). The dashed line is experimental values [85].
Separated flow over a bump
@@
Figure 6: (A) Schematic of the Gaussian bump; (B) grid structure and (C) instantaneous streamwise velocity in the $y/L=0$ plane. The velocity of the freestream is $U_{\infty}$. Note that the figure does not show the whole domain in the streamwise direction. (D) Wall pressure coefficient $C_{p}=\frac{p-p_{\infty}}{\frac{1}{2}\rho_{\infty}U^{2}_{\infty}}$ and (E) wall friction coefficient $C_{f}=\frac{\tau_{w}}{\frac{1}{2}\rho_{\infty}U^{2}_{\infty}}$, along the $y/L=0$ plane, where $U_{\infty}$, $p_{\infty}$ and $\rho_{\infty}$ are the freestream velocity, pressure and density, respectively, and $p$ and $\tau_{w}$ are averaged over time. (F) Wall pressure coefficient along the $x/L=0$ plane. (G) Mean velocity profile at $y/L=0$ at different streamwise locations. In (D–G) The solid lines denote BFM (blue), DSM-EQ (red) and VRE-EQ (yellow). White circles are experimental values[86, 87].
@@
The second validation case is a Gaussian bump designed by The Boeing Company and the University of Washington [88], as shown in Fig. 6. The test consists of a three-dimensional tapered hump at a Reynolds number of $3.6\times 10^{6}$ and a Mach number of $0.17$, based on the freestream quantities and the length of the bump along the spanwise direction, $L$. This test case has proven to be highly challenging for the RANS and LES community, with inaccurate prediction of the separation bubble and non-monotonic convergence of the solution with grid refinements[89, 90] (see Fig. 10C–D).
@@
We evaluate the performance of the BFM in predicting the mean velocity profiles, and the mean pressure and friction at the wall (characterized by the non-dimensional coefficients $C_{p}$ and $C_{f}$, respectively). The simulation consists of roughly 9 million control volumes. Similar WMLESes are conducted using DSM-EQ and VRE-EQ. The numerical findings are compared with experimental data[86]. Figure 6(D,E) illustrates the evolution of $C_{p}$ and $C_{f}$ along the centerline of the bump ($x/L$). Notably, the BFM yields more accurate predictions of the $C_{p}$ and $C_{f}$ evolution in the separation region ($x/L>0.1$) compared to DSM-EQ and VRE-EQ, which even fail to predict flow separation. The improved predictions by the BFM also extend to the presssure along the spanwise direction at the bump apex ($x/L=0$), as shown in Fig. 6F, and the mean velocity profiles, as demonstrated in Fig. 6G. The BFM predicts mean velocity with up to 3% accuracy, whereas both DSM-EQ and VRE-EQ yield a similar result that overestimate the mean velocities by more than 30%. Similarly, the BFM accurately predicts the flow topology of the separation. This is illustrated in Fig. 7, which shows a reduced window (with limits $z/L=\pm 0.13$ and $y/L\leq 0.075$) downstream of the bump. The BFM captures the counter-rotating vortices that lift away from the surface[86], in contrast to VRE-EQ, which predicts a completely different flow topology. The results for DSM-EQ (not shown) are similar to those of VRE-EQ. Nevertheless, there is still room for improvement upstream of the bump ($x/L<0$), where all models exhibit low accuracy in the prediction of $C_{f}$ (Fig. 6E).
[…]
Figure 7: Mean streamwise velocity (colormap) and in-plane streamlines at the plane $x/L=0.208$. (A) experiments[86], (B) BFM and (C) Vreman-EQ. Colors range from (dark blue) $U=-0.3U_{\infty}$ to (dark red) $U=1.4U_{\infty}$. Note that axes are not to scale.
Aircraft in landing configuration
@@
The third validation case is an aircraft in landing configuration, specifically the NASA Common Research Model High-Lift (CRM-HL). This case serves as the gold standard used by the aerospace community to assess the capabilities of different CFD methodologies [91]. The CRM-HL is a geometrically realistic aircraft that includes the bracketry associated with deployed flaps and slats, as well as a flow-through nacelle mounted on the underside of the wing (Fig. 2). The Reynolds number of the aircraft is 5.49 million based on the mean aerodynamic chord and freestream velocity, and the freestream Mach number is 0.2.
@@
The simulations are conducted using a grid with approximately 30 million control volumes. Four angles of attack are considered: $\alpha=7.05^{\circ}$, $11.29^{\circ}$, $17.05^{\circ}$ and $19.57^{\circ}$. Figure 2 contains the model geometry, an inset of the mesh, and a visualization of the wall-shear stress predicted by BFM.
Figure 8(A-C) displays the forces acting on the aircraft, characterized by the lift ($C_{L}$), drag ($C_{D}$), and pitching moment ($C_{M}$) coefficients for BFM, DSM-EQ, and VRE-EQ. The results are compared with experimental data [92]. The three plots should be interpreted as a whole, as recent studies[93] have demonstrated that –due to the complexity of the case– accurate predictions of $C_{L}$ and $C_{D}$ may result from error cancellation during the integration of aerodynamic forces. This issue is less significant in the pitching moment, which is sensitive to force distribution and hence less susceptible to error cancellation. Overall, BFM offers similar or improved accuracy in predicting the three aerodynamic coefficients, especially at high angles of attack during stall and post-stall phases. It is important to note that despite the accurate predictions of lift and drag by DSM-EQ and VRE-EQ at $\alpha=7.05^{\circ}$, these are accompanied by a poorer prediction of the pitching moment compared to BFM.
@@
Inspection of the sectional pressure coefficient ($C_{p}$) along the wing, as depicted in Fig. 8(D,E) for two spanwise locations, reaffirms that BFM yields more accurate predictions compared to DSM-EQ and VRE-EQ. The improvements are more noticeable in the flaps (last 20% of the wing chord). The results are still far from satisfactory, especially for the pitching moment, and there is clear potential for further improvements. Nonetheless, it is worth noting that BFM has never ‘seen’ an aircraft-like flow or been trained in a case resembling an airfoil or a wing. This demonstrates the ability of the BFM methodology to offer predictions that go beyond simple regression.
@@
Figure 8: (A) Lift coefficient, (B) drag polar and (C) pitching moment coefficient for the CRM-HL. The black lines denote experimental results, blue circles are for the BFM, squares are for DSM-EQ and orange squared for VRE-EQ. Pressure coefficient along the chord of the wing at (D) the 33% and (E) 55% spanwise section of the wing for $\alpha=17.05^{\circ}$. The chord location is normalized by the sectional chord-length $c_{\text{ref}}$. Empty circles represents experimental results, blue circles are for the BFM, orange and green plus symbols are for DSM-EQ and VRE-EQ, respectively.
Juncture flow in simplified aircraft
We evaluate the performance of the BFM using data from the NASA Juncture Flow Experiment[9]. This experiment involves a simplified aircraft configuration, as illustrated in Fig. 9, featuring flow separation at the juncture between the fuselage and the trailing edge of the wing[9]. The Reynolds number is calculated based on the fuselage length ($L=4839.2$ mm) and the free-stream velocity, yielding $\textit{Re}=20.8\times 10^{6}$. Additionally, the free-stream Mach number is $M=0.189$, and the angle of attack is $\alpha=5^{\circ}$. A cross-section of the grid, containing approximately 30 million points, is depicted in Fig. 9D. This analysis complements the CRM-HL case previously discussed, as the experimental campaign of the Juncture Flow contains mean velocity profiles at various locations, which were not available for the CRM-HL. Unlike integrated measurements such as lift and drag coefficients, mean velocity profiles are less susceptible to error cancellation, thus facilitating a more detailed evaluation of the models.
WMLES was conducted using both BFM and VRE-EQ. The results for the mean streamwise ($U$) and spanwise ($W$) velocity components are compared with experimental measurements at three different locations, as shown in Figure 9(A–C). The symbols in the figure represent the actual locations of the grid points. BFM yields accurate results in sections A and B, showing an improvement over VRE-EQ. BFM also improves predictions at section C. However, both models fail to capture the flow separation. This failure is not surprising, given that the thickness of the separation region is represented by only two grid points in the wall-normal direction, making accurate prediction extremely challenging. Despite this limitation, BFM still provides enhanced predictions of the mean velocity profile compared to VRE-EQ across the same grid resolution at all three locations. WMLES was also performed using DSM-EQ (not shown), which yielded slightly poorer predictions than VRE-EQ.
Figure 9: Mean velocity profile in the NASA Juncture Flow Experiment. The probes are located at (A) the upstream region of the fuselage ($x=1168.4$ mm, $z=0$ mm), (B) the wing-body junction ($x=2747.6$ mm; $y=239.1$ mm), and (C) the wing-body junction close to the trailing edge ($x=2922.6$ mm; $y=239.1$ mm). The approximate location are depicted in the sketch at the top. In all the plots, the symbols denote: BFM (blue circles) and VRE-EQ (yellow triangles). Closed symbols corresponds to $U/U_{\infty}$ and open symbols to $W/U_{\infty}$, where $U_{\infty}$ is the freestream velocity. The lines correspond to experimental values[9]. $\delta$ is the local boundary layer thickness [26]. (D) Cross-section of the grid at the wing.
Discussion
Future opportunities of BFM for WMLES